Mary studies 8000 hours per year. David studies 260 hours per year. how many years will david study to have 8000 hours?
Answer by RedemptiveMath(80)

For every year that goes by David studies 260 hours. That is, in one year he studies 260 hours (total), in two years he studies for 520 hours (total), in three years he studies for 780 hours (total), and so on and so forth. This question wants to know at what point (in years) David will reach 8000 hours (total) of study. All we need to do is divide 8000 hours by 260 hours per year to get how many years it will take him to reach the mark:

8000/260 = 400/13 or 30 and 10/13 or approx. 30.769 years.

It is dependent on your teacher what form you write the answer in. Again, every year that goes by David studies for 260 hours. After about 30.8 years, he would have studied for 8000 hours total. You would know to divide this way because you are given a total amount of hours, 8000, and what we call a rate, or 260 hours per year. We usually want to divide the totals by the rates in order to get missing information. Think of it like this: We have two totals, total hours (8000) and total years (we don't know), and we have a rate (260 hours per year) that links the two totals together. Rates link two totals together in order to make things simple. Whenever you want to find a rate, just divide the two totals together with respect to what we are working with. For example, if we didn't have our rate of 260 hours per year and we did have the total amount of years (400/13), we would divide the total amount of hours by the total amount of years to get the rate (hence the rate says "hours per year"). 8000 divided by 400/13 = 260.